Trying to help my daughter...
kali mixes together 9 gal of brand a fruit drink and 6 gal of brand b fruit drink which contains 5% fruit juice. find the percent of fruit juice in brand a if the mixture contained 11% fruit juice.
Thanks

Answer by stanbon(75887)

kali mixes together 9 gal of brand a fruit drink and 6 gal of brand b fruit drink which contains 5% fruit juice. find the percent of fruit juice in brand a if the mixture contained 11% fruit juice.
-------
Equation:
concentrate + concentrate = concentrate
Let the unkown percent in "b" be x.
-----------------------------------------
0.05*6 + x*9 = 0.11(6+9)
---
Multiply thru by 100 to get:
5*6 + 900x = 11*15
---
30 + 900x = 165
900x = 135
x = 0.15
----
Ans: Percent of fruit juice in Brand b is 15%
=================
Cheers,
Stan H.
======================